OpenStudy (anonymous):

5x+6y=7 -2x+y=21 substitution method

OpenStudy (anonymous):

y = 2x + 21 plug this into 5x + 6y = 7 to give an equation in x solve and sub this value in one of the original equatiuos to get y

OpenStudy (anonymous):

The same method again. take the 2nd one and give a result for y. It will go like this: y=2x+21. Now go at the 1st one and when you see y put 2x+21. It will go like this: 5x+6(2x+21)=7...... you will found x and then go at 2nd again and put the number that you found. Use the same method for all.

OpenStudy (anonymous):

so x=-7?

OpenStudy (anonymous):

y=126?

OpenStudy (anonymous):

x=-7,right. So,now you will go at y=2x+21 and whenever you see x you will put -7. So we have : y= 2(-7)+21=-14+21=7 A(-7,7).

OpenStudy (anonymous):

I think I get it... I just do not know why I am having a hard time with this..

OpenStudy (anonymous):

I have another, I am going to try this one and let me know if I am correct.. if you do not care please

OpenStudy (anonymous):

It's pretty easy. That's one of the 3 types that you can solution this type of exercices. The 2nd type it's to multiply the hole equation with a number that you can add from the other equation for have 0 on x or y. Then you can easy find the other. When I say it it's harder to understand it. Focus more at school :)

OpenStudy (anonymous):

8x+5y=-4 -2x+y=10 8x+5(2x+10)=-4 x=-3 2(-3)+10=4+10=14 (4,-4)?

OpenStudy (anonymous):

x=-3,that's right. Now you are going at -2x+y=10 and whenever you see x you will put -3. so we have : (-2)(-3)+y=10 => y=10-6 => y=4. A(-3,4). Good job,you are going better. :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem my friend.

OpenStudy (anonymous):

This one is a little different, can you help with this one please? 7x-6y=-35 5x+48=y

OpenStudy (anonymous):

7x-6(5x+48)=-35 x=-11 5(-11)+48=-7 y=-7

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

Omg... Thank you!!